OpenStudy (anonymous):

Andrew plays on a basketball team. In two games he scored 2/5 of the total number of points his team scored. His team scored 55 points in the first game and 35 points in the second game. What was the number of points Andrew scored in these two games? A. 18 B. 20 C. 36 D. 90

OpenStudy (anonymous):

please give me an answer

OpenStudy (anonymous):

\[\frac{2 (55+35)}{5}=36 \]( 2/5 )( 55 + 35 ) = 36

OpenStudy (anonymous):

oh so its C

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

You are welcome.